# WISE: A World Knowledge-Informed Semantic Evaluation for Text-to-Image Generation

### v1

@@ 3.1 WISE construction
Unify MLLM
@@
show-o-demo 0.28 0.36 0.40 0.23 0.33 0.22 0.30
show-o-demo-512 0.28 0.40 0.48 0.30 0.46 0.30 0.35
@@
Cultural Common Sense. The Cultural Common Sense domain of WISE aims to evaluate models’ ability to understand and apply culturally specific knowledge in image generation, reflecting a crucial aspect of real-world understanding. A model should not only understand the visual features of an object (e.g., shape, size, color of a turkey) but also align with real-world cultural knowledge, such as the cultural significance of a turkey as a Thanksgiving food. A model lacking this deeper understanding would be an incomplete image generator, only capable of grasping the shallow alignment between nouns and visual forms, without understanding the object’s cultural role in the real world. This section encompasses a diverse range of topics, categorized into 10 fine-grained sub-domains, including festivals, sports, religion, crafts, architecture, animals, plants, art, celebrities, and daily life. These categories collectively cover a broad spectrum of human cultural experiences and knowledge. For example, prompts may involve generating images related to traditional festival customs, characteristic national crafts, iconic landmarks, animals and plants with specific cultural significance, common knowledge from daily life, or events and objects associated with famous figures.
Spatio-temporal Reasoning. The Spatio-temporal Reasoning domain in WISE is structured around two key dimensions: Temporal Reasoning and Spatial Reasoning. Temporal Reasoning is divided into Horizontal Temporal, which assesses the understanding of relative temporal relationships between events or objects (e.g., “The Statue of Liberty at 10 PM Dubai time.”), and Vertical Temporal, which evaluates the comprehension of absolute temporal relationships, involving specific points in time (e.g., morning, noon, evening, seasons, specific years, or centuries). Spatial Reasoning is divided into three subcategories: Different Views, which tests the understanding of different viewpoints, including top view, bottom view, cross-section, side view, mirror image, and the effects of perspective; Geographical Relationships, which assesses the understanding of spatial relationships between cities, countries, continents, and other geographical entities; and Relative Position, which focuses on the understanding of the spatial arrangement of objects relative to each other within a scene.
Natural Science. Finally, the WISE benchmark includes a Natural Science domain, designed to assess whether models can not only comprehend domain-specific scientific knowledge but also apply this understanding to reason about complex scientific scenarios and generate accurate, scientifically consistent images. Generative models, at their core, are designed to model the real world. We aim to evaluate whether these models’ understanding goes beyond mere visual replication and encompasses the complex scientific knowledge underlying real-world phenomena. For example, a model should not only be able to generate images of water, ice, and steam, but also understand the underlying thermodynamic principles (e.g., freezing, evaporation, condensation) that govern the transitions between these states. This deeper understanding would demonstrate that the model is not simply mimicking visual appearances but is also incorporating fundamental scientific facts, such as how temperature influences the state of water.
This domain moves beyond general knowledge and delves into specialized concepts in biology, physics, and chemistry. The Natural Science domain comprehensively evaluates models’ understanding of biological processes and states across different life cycles (e.g., metamorphosis, seasonal changes), time periods (e.g., diurnal and nocturnal variations), and environmental conditions (e.g., responses to stimuli, resource availability), as well as complex animal behaviors (e.g., predation, defense, migration); fundamental physical principles and phenomena including mechanics (gravity, buoyancy, pressure, tension), thermodynamics (phase transitions: evaporation, condensation, freezing, melting, sublimation), optics (refraction, reflection, magnification, dispersion), and material properties (e.g., conductivity, elasticity); and combustion, metal corrosion, solution chemistry (acid-base reactions, precipitation reactions, displacement reactions), and characteristic chemical properties (e.g., Tyndall effect, allotropy, ion-specific solution colors).
3.2 Evaluation Framework
Evaluation method In recent years, Multimodal Large Language Models (MLLMs) have achieved remarkable advancements in image understanding, leveraging the world knowledge embedded within large language models to attain strong perception and comprehension capabilities. To mitigate the limitations of CLIP’s inherent world knowledge constraints, we employ GPT-4o, a powerful MLLM, as the judge to evaluate the performance of T2I models in this work. We adopted a meticulously designed and rigorous scoring mechanism, with details illustrated in Appendix.
Table 2: Normalized WiScore results on rewritten prompts. These prompts were simplified from the original WISE benchmark using GPT-4o (e.g., “The plant often gifted on Mother’s Day” to ”Carnation”).
Model Cultural Time Space Biology Physics Chemistry Overall
Dedicated T2I
FLUX.1-dev 0.75 0.70 0.76 0.69 0.71 0.68 0.73
FLUX.1-schnell 0.63 0.58 0.67 0.58 0.58 0.44 0.60
PixArt-Alpha 0.66 0.64 0.55 0.58 0.64 0.62 0.63
playground-v2.5 0.78 0.72 0.63 0.69 0.67 0.60 0.71
SD-v1-5 0.59 0.50 0.41 0.47 0.44 0.36 0.50
SD-2-1 0.63 0.61 0.44 0.50 0.49 0.41 0.55
SD-XL-base-0.9 0.68 0.71 0.59 0.61 0.67 0.55 0.65
SD-3-medium 0.76 0.65 0.68 0.59 0.67 0.59 0.69
SD-3.5-medium 0.73 0.69 0.67 0.68 0.67 0.60 0.69
SD-3.5-large 0.78 0.69 0.68 0.64 0.70 0.64 0.72
Unify MLLM
Emu3 0.70 0.62 0.60 0.59 0.56 0.52 0.63
Janus-1.3B 0.40 0.48 0.49 0.54 0.53 0.44 0.46
JanusFlow-1.3B 0.39 0.43 0.38 0.57 0.44 0.41 0.42
Janus-Pro-1B 0.60 0.59 0.59 0.66 0.63 0.58 0.60
Janus-Pro-7B 0.75 0.66 0.70 0.71 0.73 0.59 0.71
Orthus-7B-base 0.19 0.23 0.20 0.24 0.21 0.21 0.21
Orthus-7B-instruct 0.55 0.47 0.48 0.46 0.45 0.42 0.50
show-o-demo 0.61 0.56 0.55 0.54 0.53 0.56 0.57
show-o-demo-512 0.64 0.62 0.68 0.63 0.69 0.59 0.64
vila-u-7b-256 0.54 0.51 0.49 0.57 0.56 0.58 0.54
Evaluation Metrics We propose a multi-faceted evaluation protocol to rigorously assess the quality of generated images, focusing on four key aspects: Consistency, Realism, Aesthetic Quality, and a composite metric, WiScore, shown in Figure 4. Consistency evaluates the accuracy and completeness with which the generated image reflects the user’s prompt, capturing all key elements and nuances. Realism assesses the realism of the image, considering adherence to physical laws, accurate material representation, and coherent spatial relationships, determining how closely the image resembles a real photograph. Aesthetic Quality measures the overall artistic appeal and visual quality of the image, encompassing aspects such as composition, color harmony, and artistic style. WiScore, the central metric, emphasizes the accuracy of the depicted objects or entities within the generated image, directly reflecting our benchmark’s focus on world knowledge utilization. It is calculated as a weighted value of the other three metrics:
[…]
The weighting prioritizes Consistency, reflecting the importance of accurately representing the prompt’s intended objects and their relationships, while also incorporating Realism and Aesthetic Quality to ensure overall image quality. A higher WiScore indicates superior performance in accurately depicting objects and concepts based on world knowledge. Detailed Scoring Criteria in Figure 5. To facilitate comparison, in Table 1 and Table 2, we report the average WiScore per image, normalized by dividing by 2. We provide complete results on each score in Appendix.
4 Evaluation Results
4.1 Experiment settings
We evaluate 20 T2I models, including 10 dedicated text-to-image models and 10 unified multimodal models. The dedicated T2I models are: stable-diffusion-v1-5 [33], stable-diffusion-2-1 [33], stable-diffusion-xl-base-0.9 [29], stable-diffusion-3-medium [5], stable-diffusion-3.5-medium [5], stable-diffusion-3.5-large [5], playground-v2.5-1024px-aesthetic [20], PixArt-XL-2-1024-MS [1], FLUX.1-dev [18], and FLUX.1-schnell [18]. The unified multimodal models are: Janus-Pro-7B [4], Janus-Pro-1B [4], JanusFlow-1.3B [25], Janus-1.3B [44], show-o-demo [49], show-o-demo-512 [49], Orthus-7B-base [17], Orthus-7B-instruct [17], vila-u-7b-256 [48], and Emu3 [43]. For image generation, we use the official default configurations of each model and fix the random seed to ensure reproducibility. We utilize GPT-4o-2024-05-13 as the evaluation model. All experiments are conducted on eight NVIDIA A800 GPUs. For short, we use SD to denote stable-diffusion, playground-v2.5 for playground-v2.5-1024px-aesthetic, and PixArt-Alpha for PixArt-XL-2-1024-MS.
4.2 Overall results
As the results in Table 1 show, generating images that accurately and comprehensively integrate world knowledge remains a substantial challenge across all models, encompassing both dedicated text-to-image models and unified multimodal models. This highlights a critical area for improvement in the T2I field, suggesting a significant deficiency in current approaches when it comes to complex semantic understanding and the integration of world knowledge for image generation. Furthermore, a clear distinction arises between dedicated T2I models and unified multimodal models. Notably, FLUX.1-dev, leveraging a CLIP [30] text encoder and T5 [31], achieves an overall WiScore of 0.50. This underscores the potential of combining CLIP’s visual language alignment with T5’s powerful language representation for effective knowledge grounding, yet also underscores the persistent challenges that remain, as even this architecture falls short of achieving truly high scores, highlighting the need for even more advanced approaches.
Understanding-Generation gap of unified multimodal models. As results in Table 1 indicate, dedicated T2I models generally outperform unified multimodal models in terms of overall WiScore. This suggests that despite the robust text and image understanding capabilities afforded by training on large-scale image-text pairs and massive textual datasets, unified multimodal models have yet to fully leverage their inherent strengths for effective knowledge integration in the image generation process. This gap signifies that the capability to understand and reason about the world does not automatically translate into the ability to visually represent that knowledge with sufficient fidelity and accuracy in current unified multimodal models.
Performance of different categories. Furthermore, while performance varied across all categories, cultural prompts generally elicited the most accurate and coherent images, suggesting a relative strength in handling common knowledge and readily available cultural concepts. In contrast, prompts about natural science and space-time consistently yielded lower scores, demonstrating a clear weakness in accurately representing these more complex and specialized domains. This difference in performance likely stems from a combination of factors. The superior handling of cultural prompts likely reflects the prevalence of cultural knowledge and common-sense information within the training datasets used to build these models. Conversely, accurate visual representation of science and space-time concepts often necessitates more sophisticated reasoning abilities and the integration of less frequently encountered information. Additionally, the relative scarcity of high-quality, accurately labeled training data for these technical domains may further hinder the models’ ability to effectively translate prompts into corresponding visual representations.
Effect of language model parameters. Janus-Pro-7B exhibits a noticeable performance increase compared to Janus-Pro-1B, highlighting the role of language model scale. It underscores that larger language models, with their increased parameter count, possess a greater capacity for encoding world knowledge and a stronger ability to understand and utilize complex textual information for image generation.
@@ 4.2 Overall results
We conducted an additional experiment to further highlight the limitations of current T2I models in handling world knowledge. We used GPT-4o-2024-05-13 to rewrite the prompts in our WISE benchmark, transforming them from complex, knowledge-demanding prompts into direct prompts. For example, the original prompt “The plant often gifted on Mother’s Day” was rewritten as “Carnation.” The specific instructions for prompt rewriting are detailed in Appendix. The results of this experiment, using the same WiScore evaluation, are presented in Table 2. Nearly all models exhibit a significant performance increase when evaluated with the rewritten prompts. Notably, dedicated T2I models like FLUX.1-dev and SD-3.5-large, as well as the unified multimodal model Janus-Pro-7B, show substantial improvements in their overall WiScore. However, even with this simplification, the scores, while improved, still do not reach a level that would indicate a complete and satisfactory understanding of world knowledge across all categories. This demonstrates the instability and limitations of relying solely on LLMs to rewrite prompts for evaluating world knowledge in T2I models. The substantial performance variation highlights the need for future research to focus on improvements in model training methodologies, rather than relying on prompt engineering alone.
@@ 5 Conclusion
To evaluate the ability of current T2I models to generate images that goes beyond simple word-pixel mapping, we introduce WISE, a novel benchmark with 1000 prompts spanning diverse domains, designed to challenge T2I models’ ability to generate images grounded in common world knowledge. Our evaluation of 20 T2I models, including both dedicated T2I models and unified multimodal models, reveals significant shortcomings in their capacity to effectively leverage world knowledge during image generation. Even unified multimodal models, despite their strong language understanding capabilities, do not fully translate this advantage into superior image generation performance in complex, knowledge-demanding scenarios.
@@ 6 Limitations
While our work introduces a novel benchmark, WISE, for evaluating the world knowledge and semantic understanding capabilities of T2I models, we acknowledge several limitations. Our benchmark categorizes prompts into broad domains, but due to the interconnected nature of knowledge, some prompts may inherently span multiple categories (e.g., “the impact of climate change on polar bear habitats” could fall under both natural science and spatio-temporal reasoning), potentially introducing ambiguity in cross-category analysis. Furthermore, while WISE covers a range of topics, it represents a sample of knowledge domains and cannot encompass all aspects of world knowledge, which is also constantly evolving. Additionally, some models were not publicly available or did not provide APIs at the time of our work’s deadline, precluding their evaluation in this study.
Acknowledgements. We are grateful to Jiaxiu Liu, Chaoran Feng, Fanqing Meng and Hongcheng Gao for helpful discussion and drawing.
@@ Appendix A WISE Category Descriptions
WISE (likely an acronym for a knowledge evaluation framework, not specified in the provided context) encompasses a broad spectrum of knowledge categories, categorized under three main domains: Cultural Common Sense, Spatio-Temporal Reasoning, and Natural Science. Each domain is further divided into specific subcategories to assess different facets of understanding.
@@ A.3.3 Chemistry (100 prompts)
This structured categorization of WISE provides a comprehensive framework for evaluating a model’s knowledge across diverse domains, ranging from cultural understanding to scientific principles.
Figure 7: We utilize GPT-4 rewrite the prompts in our WISE benchmark, transforming them from complex, knowledge-demanding prompts into direct prompts. Above is the instruction we provided to GPT-4o.
@@
show-o-demo 172.00 109.00 104.00 32.00 51.00 24.00 111.53
show-o-demo-512 156.00 107.00 118.00 36.00 74.00 37.00 110.66
@@
show-o-demo 282.00 132.00 103.00 74.00 96.00 80.00 173.54
show-o-demo-512 372.00 188.00 149.00 117.00 131.00 109.00 235.71
@@
show-o-demo 395.00 170.00 131.00 98.00 113.00 104.00 235.31
show-o-demo-512 426.00 207.00 166.00 121.00 135.00 121.00 264.75
@@
Table 6: Consistency scores of different models on rewritten prompts. These prompts were simplified from the original WISE benchmark using GPT-4o (e.g., “The plant often gifted on Mother’s Day” to ”Carnation”).
@@
show-o-demo 518.00 191.00 152.00 107.00 102.00 115.00 291.71
show-o-demo-512 524.00 195.00 185.00 123.00 133.00 114.00 303.77
@@
Table 7: Realism scores of different models on rewritten prompts. These prompts were simplified from the original WISE benchmark using GPT-4o (e.g., “The plant often gifted on Mother’s Day” to ”Carnation”).
@@
show-o-demo 392.00 167.00 125.00 102.00 114.00 94.00 232.31
show-o-demo-512 458.00 231.00 164.00 132.00 149.00 119.00 283.59
@@
Table 8: Aesthetic Quality scores of different models on rewritten prompts. These prompts were simplified from the original WISE benchmark using GPT-4o (e.g., “The plant often gifted on Mother’s Day” to ”Carnation”).
@@
show-o-demo 496.00 205.00 149.00 119.00 128.00 124.00 289.55
show-o-demo-512 556.00 254.00 182.00 143.00 154.00 136.00 332.32

### v2

@@ 4.2 Main Results
Unified MLLM
GPT4o 0.81 0.71 0.89 0.83 0.79 0.74 0.80
MetaQuery-XL 0.56 0.55 0.62 0.49 0.63 0.41 0.55
Liquid 0.34 0.45 0.48 0.41 0.45 0.27 0.39
@@
show-o 0.28 0.36 0.40 0.23 0.33 0.22 0.30
show-o-512 0.28 0.40 0.48 0.30 0.46 0.30 0.35
@@
Understanding-Generation Gap in T2I Tasks.Table 1 presents the WiScore of various models on the WISE benchmark, designed to evaluate the extent to which models leverage world knowledge in image generation. Overall, the results indicate that the GPT4o model significantly outperforms all others across all categories and the overall score. This suggests its superior capabilities in utilizing complex semantic understanding (implicit understanding) and world knowledge (intrinsic knowledge matching) for image generation. The remaining open-source models demonstrate comparatively lower performance, with Unified MLLMs generally lagging behind Dedicated T2I models. This highlights that while Unified MLLMs are trained on large-scale image-text pairs and vast textual datasets, thereby acquiring substantial text and image understanding capabilities, effectively integrating this knowledge into the image generation process remains a significant challenge. Despite the overall underperformance of Unified MLLMs, the MetaQuery-XL model exhibits a prominent performance, achieving the best score among the open-source models, showcasing the potential of unified models to apply world knowledge for image generation.
Cross-Dimensional Performance Statical Analysis. This part systematically examines the performance disparities between the two model categories across various cognitive dimensions, employing a suite of statistical methods including paired-sample t-tests, Wilcoxon signed-rank tests, and effect size analysis. The paired t-test results indicate that T2I models demonstrate a significantly higher average score in the first three dimensions (Cultural/Time/Space) ($0.56\pm 0.08$) compared to the latter three (Biology/Physics/Chemistry) ($0.43\pm 0.07$), $t(9)=5.47$, $p<0.001$, with an effect size of Cohen’s $d=1.73$. This suggests a highly statistically and practically significant difference. To verify the robustness of these findings, a non-parametric Wilcoxon signed-rank test was conducted, yielding consistent conclusions ($W=55$, $p=0.002$). Similarly, the Unified MLLM model group also exhibited a characteristic pattern where the first three dimensions ($0.80\pm 0.19$) significantly outperformed the latter three ($0.79\pm 0.19$) ($t(10)=3.76$, $p=0.003$, $d=1.09$; Wilcoxon $W=66$, $p=0.005$). A Mann-Whitney U test revealed that although the magnitude of inter-dimensional differences was significantly larger for T2I models ($\Delta=0.13$) compared to Unified MLLMs ($\Delta=0.01$) ($U=32$, $p=0.057$), all models displayed an advantage in cultural tasks. This pattern of differences may stem from: 1) the Cultural/Time/Space dimensions relying more heavily on visual representation capabilities, a known strength of T2I models; and 2) scientific tasks requiring stronger cross-modal reasoning abilities, in which MLLM models (such as GPT4o) exhibit more balanced performance ($\Delta=0.01$). Notably, within the SD series, from v1.5 to 3.5-large, the dimensional difference decreased from 0.08 to 0.05; likewise, in the Janus series, from 1.3B to Pro-7B, the difference diminished from 0.04 to 0.03. This suggests that newer generation models are improving in their ability to balance performance across different dimensions.
@@
Figure 5: WiScore on rewritten prompts of different models. These prompts were simplified from the original WISE benchmark using GPT-4o (e.g., "The plant often gifted on Mother’s Day" to "Carnation"). Green ball indicates score increase after rewriting; red ball indicates score decrease.
We conducted an additional experiment to further highlight the limitations of current T2I models in processing world knowledge and explore possible solutions. We rewrote the prompts in the WISE benchmark using GPT-4o-2024-05-13, transforming them from complex, knowledge-requiring prompts to straightforward prompts. For example, the original prompt "Common plants for Mother’s Day" was rewritten to "Carnation". The detailed instructions for prompt rewriting are detailed in Appendix F. The same WiScore evaluation method was used for this experiment, and the results are presented in  Figure 5.
Surprisingly, nearly all models demonstrated significant performance improvements when evaluated with the rewritten prompts. The average WiScore increase across all Dedicated T2I models was +0.22, while Unified MLLMs saw an even larger average increase of +0.27. This suggests that a significant portion of the difficulty in the original WISE benchmark stemmed from the models’ inability to effectively parse and contextualize complex, knowledge-intensive prompts. By simplifying the prompts, we effectively removed this bottleneck, allowing the models to better showcase their generative capabilities. Notably, the models with the largest overall improvements tended to be those with lower initial scores on the original benchmark, indicating that prompt complexity had a disproportionately negative impact on their performance. For example, Janus-Pro-1B demonstrated an overall WiScore improvement of +0.34 after rewriting, highlighting the sensitivity of weak models to prompt complexity.
However, even with this simplification, the scores of both T2I and unified models, while improved, did not reach a level indicative of a complete and satisfactory generation of world knowledge across all categories. Furthermore, it highlights the core issue of lower-quality generation results in current unified models. Even MetaQuery-XL experienced a decrease in its score for the Physics category. This demonstrates that solely relying on enhancing the model’s ability to interpret world knowledge remains constrained by the bottleneck of generative capacity. Concurrently, the substantial performance gap underscores the need for future research to focus on improving model training methods and architectures, rather than solely relying on tricks aimed at enhancing the understanding level, such as prompt engineering.
@@ 5 Conclusion
To evaluate the ability of current T2I and unified models to generate images beyond simple bag-of-words mappings, we introduced WISE, a new benchmark consisting of 1,000 questions covering different domains designed to challenge the ability of T2I models to generate images based on common world knowledge. We evaluated 22 generative models (including dedicated T2I models and unified multimodal models) and found that these models have significant deficiencies in their ability to effectively utilize world knowledge in the image generation process, especially the T2I model. Even for the unified multimodal model, despite its strong language understanding capabilities, most open-source models are unable to fully convert this advantage into excellent image generation performance in complex and knowledge-intensive scenarios. From this point of view, we also pointed out through some analysis that the key to the potential problem lies in the improvement of image generation quality itself. Therefore, this paper points out the deficiencies of existing generative models from a benchmark perspective and points out an effective path for subsequent methods.
@@ 6 Limitations
While our work introduces a novel benchmark, WISE, for evaluating the complex semantic understanding (implicit understanding) and world knowledge (intrinsic knowledge matching) capabilities of dedicated T2I and unified models, we acknowledge several limitations. Our benchmark categorizes prompts into broad domains, but due to the interconnected nature of knowledge, some prompts may inherently span multiple categories (e.g., "the impact of climate change on polar bear habitats" could fall under both natural science and spatio-temporal reasoning), potentially introducing ambiguity in cross-category analysis. Furthermore, while WISE covers a range of topics, it represents a sample of knowledge domains and cannot encompass all aspects of world knowledge, which is also constantly evolving. Additionally, some models were not publicly available or did not provide APIs at the time of our work’s deadline, precluding their evaluation in this study.
Acknowledgements. We are grateful to Xichen Pan, Jiaxiu Liu, Fanqing Meng, and Hongcheng Gao for helpful discussions and drawing.
@@ Appendix A WISE Category Descriptions
WISE encompasses a broad spectrum of knowledge categories, categorized under three main domains: Cultural Common Sense, Spatio-Temporal Reasoning, and Natural Science. Each domain is further divided into specific subcategories to assess different facets of understanding.
@@ A.3.3 Chemistry (100 prompts)
Appendix B Selection of WiScore Hyperparameters
Human Evaluation.We randomly sampled 100 pictures generated by FLUX.1-dev. We require paper authors to rate each image according to the same standard Figure 7, and regard the sum of the final values as the score.
Table 2: Evaluation metrics comparison
Metric Human 4o
Consistency ($C$) 96 84
Realism ($R$) 112 153
Aesthetic Quality ($A$) 143 159
Given metrics from Table 2, solve:
$\min_{\mathbf{w}}\left\|\begin{bmatrix}12&-41&-16\end{bmatrix}\mathbf{w}\right\|^{2}\quad\text{s.t.}\quad\begin{cases}w_{1}+w_{2}+w_{3}=1\\ w_{i}\geq 0\ \forall i\end{cases}$ (2)
The KKT conditions yield:
$0\displaystyle 12w_{1}-41w_{2}-16w_{3}=$ 0 (3)
$\mathbf{w}=\left(\frac{41-25w_{3}}{53},\ \frac{12-28w_{3}}{53},\ w_{3}\right)$ (4)
Parameter $w_{3}$ spans:
[…]
Table 3: Parameter grid ranges ($\Delta=0.1$)
Parameter Search Range
$w_{1}$ [0.5, 0.8]
$w_{2}$ [0.1, 0.3]
$w_{3}$ [0.1, 0.4]
[…]
Figure 6: 3D visualization of grid search results showing $(0.7,0.2,0.1)$ as the minimal point
The optimal weights were found at:
$\mathbf{w}^{*}=(0.7,0.2,0.1)$ (6)
achieving the minimal squared difference of 1.8.
Appendix C Scoring Criteria and Results of Other Metrics
@@
show-o 172.00 109.00 104.00 32.00 51.00 24.00 111.53
show-o-512 156.00 107.00 118.00 36.00 74.00 37.00 110.66
@@
show-o 282.00 132.00 103.00 74.00 96.00 80.00 173.54
show-o-512 372.00 188.00 149.00 117.00 131.00 109.00 235.71
@@
show-o 395.00 170.00 131.00 98.00 113.00 104.00 235.31
show-o-512 426.00 207.00 166.00 121.00 135.00 121.00 264.75
@@
Table 7: Consistency scores of different models on rewritten prompts.
@@
show-o 518.00 191.00 152.00 107.00 102.00 115.00 291.71
show-o-512 524.00 195.00 185.00 123.00 133.00 114.00 303.77
@@
Table 8: Realism scores of different models on rewritten prompts.
@@
show-o 392.00 167.00 125.00 102.00 114.00 94.00 232.31
show-o-512 458.00 231.00 164.00 132.00 149.00 119.00 283.59
@@
Table 9: Aesthetic Quality scores of different models on rewritten prompts.
@@
show-o 496.00 205.00 149.00 119.00 128.00 124.00 289.55
show-o-512 556.00 254.00 182.00 143.00 154.00 136.00 332.32
@@
Appendix D GPT4o’s Assessment Instruction
Figure 8 presents the instructions we input to GPT4o for evaluation.
Appendix E Comparison of Human Assessment with Various Metrics
We randomly sampled 100 images generated by FLUX.1-dev, and reproduced CLIPscore and VQAscore according to the official settings, and calculated the scores in categories. The results are shown in Figure 4. It can be seen that the WiScore fits human scores best.
Appendix F Using GPT4-o Rewrite Prompts on WISE
Figure 9 presents the instructions we input to GPT4o to rewrite prompts.
Figure 9: We utilize GPT-4o to rewrite the prompts in our WISE benchmark, transforming them from complex, knowledge-demanding prompts into direct prompts. Above is the instruction we provided to GPT-4o.
Table 10: WiScore on rewritten prompts of different models. These prompts were simplified from the original WISE benchmark using GPT-4o (e.g., "The plant often gifted on Mother’s Day" to "Carnation"). Green bold indicates score increase after rewriting; red bold indicates score decrease.
Model Cultural Time Space Biology Physics Chemistry Overall
Dedicated T2I
FLUX.1-dev 0.75 +0.27 0.70 +0.12 0.76 +0.14 0.69 +0.27 0.71 +0.20 0.68 +0.33 0.73 +0.23
FLUX.1-schnell 0.63 +0.24 0.58 +0.14 0.67 +0.17 0.58 +0.27 0.58 +0.14 0.44 +0.18 0.60 +0.20
PixArt-Alpha 0.66 +0.21 0.64 +0.14 0.55 +0.07 0.58 +0.09 0.64 +0.08 0.62 +0.28 0.63 +0.16
playground-v2.5 0.78 +0.29 0.72 +0.14 0.63 +0.08 0.69 +0.26 0.67 +0.19 0.60 +0.27 0.71 +0.22
SD-v1-5 0.59 +0.25 0.50 +0.15 0.41 +0.09 0.47 +0.19 0.44 +0.15 0.36 +0.15 0.50 +0.18
SD-2-1 0.63 +0.33 0.61 +0.23 0.44 +0.09 0.50 +0.17 0.49 +0.15 0.41 +0.20 0.55 +0.23
SD-XL-base-0.9 0.68 +0.25 0.71 +0.23 0.59 +0.12 0.61 +0.17 0.67 +0.22 0.55 +0.28 0.65 +0.22
SD-3-medium 0.76 +0.34 0.65 +0.21 0.68 +0.20 0.59 +0.20 0.67 +0.20 0.59 +0.30 0.69 +0.27
SD-3.5-medium 0.73 +0.30 0.69 +0.19 0.67 +0.15 0.68 +0.27 0.67 +0.14 0.60 +0.27 0.69 +0.24
SD-3.5-large 0.78 +0.34 0.69 +0.19 0.68 +0.10 0.64 +0.20 0.70 +0.18 0.64 +0.33 0.72 +0.26
Unified MLLM
MetaQuery-XL 0.81 +0.25 0.73 +0.18 0.69 +0.07 0.67 +0.18 0.59 -0.04 0.58 +0.17 0.72 +0.17
Emu3 0.70 +0.36 0.62 +0.17 0.60 +0.12 0.59 +0.18 0.56 +0.11 0.52 +0.25 0.63 +0.24
Janus-1.3B 0.40 +0.24 0.48 +0.22 0.49 +0.14 0.54 +0.26 0.53 +0.23 0.44 +0.30 0.46 +0.23
JanusFlow-1.3B 0.39 +0.26 0.43 +0.17 0.38 +0.10 0.57 +0.37 0.44 +0.25 0.41 +0.30 0.42 +0.24
Janus-Pro-1B 0.60 +0.40 0.59 +0.31 0.59 +0.14 0.66 +0.42 0.63 +0.31 0.58 +0.42 0.60 +0.34
Janus-Pro-7B 0.75 +0.45 0.66 +0.29 0.70 +0.21 0.71 +0.35 0.73 +0.31 0.59 +0.33 0.71 +0.36
Orthus-7B-instruct 0.55 +0.32 0.47 +0.16 0.48 +0.10 0.46 +0.18 0.45 +0.14 0.42 +0.22 0.50 +0.23
show-o 0.61 +0.33 0.56 +0.20 0.55 +0.15 0.54 +0.31 0.53 +0.20 0.56 +0.34 0.57 +0.27
show-o-512 0.64 +0.36 0.62 +0.22 0.68 +0.20 0.63 +0.33 0.69 +0.23 0.59 +0.29 0.64 +0.29
vila-u-7b-256 0.54 +0.28 0.51 +0.18 0.49 +0.12 0.57 +0.22 0.56 +0.17 0.58 +0.35 0.54 +0.23